OpenStudy (anonymous):

please help Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1)

jimthompson5910 (jim_thompson5910):

solve the first equation for y 2x + y = 3 y = 3 - 2x then plug this into the second equation -3x + 4y = 1 -3x + 4(3 - 2x) = 1 .. replace y with 3-2x -3x + 12 - 8x = 1 solve for x, then use this to solve for y

OpenStudy (anonymous):

so x=1?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now plug that back into y = 3 - 2x to find y

OpenStudy (anonymous):

y=3+2x?

jimthompson5910 (jim_thompson5910):

y = 3 - 2x

jimthompson5910 (jim_thompson5910):

it's minus, not plus

OpenStudy (anonymous):

so the answer is? (1,1)

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what about this one? What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 2 -2 1 -1

jimthompson5910 (jim_thompson5910):

first solve for x in x - 2y = -5 to get x - 2y = -5 x = -5+2y now plug this into the first equation 2x + 3y = 4 2(-5+2y) + 3y = 4 -10+4y + 3y = 4 solve for y, then use this to solve for x

OpenStudy (anonymous):

y=2

jimthompson5910 (jim_thompson5910):

good, then use that to find x

OpenStudy (anonymous):

how do I do that?

jimthompson5910 (jim_thompson5910):

well remember we isolated x and got x = -5+2y

jimthompson5910 (jim_thompson5910):

so if y = 2, then x = -5+2y x = -5+2*2 ... replace y with 2 (since y = 2) x = -5 + 4 x = -1

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome, keep practicing this and you'll definitely be a pro in no time

OpenStudy (anonymous):

alright I will!